# Direct-to-Cell: A First Look into Starlink’s Direct Satellite-to-Device Radio Access Network through Crowdsourced Measurements

### v7

@@ I Introduction
Following technological developments in fixed broadband connectivity, satellite-based cellular communications, known as Direct Satellite-to-Device (DS2D), have recently emerged as a promising solution to provide seamless connectivity directly to standard unmodified smartphones. Starlink, the satellite communications division of SpaceX, has emerged as the global leader in DS2D communications by rapidly deploying a non-3GPP-NTN-compliant solution that leverages terrestrial International Mobile Telecommunications (IMT) spectrum. In partnership with T-Mobile and using its spectrum, Starlink has been the first company to launch large-scale beta testing of a Supplemental Coverage from Space (SCS) service. The system is initially limited to Short Message Service (SMS), with support for voice and data services planned by late 2025.
@@ II-A Technological Evolution and Emerging Ecosystem
Starlink has pioneered DS2D deployments with a strategy focused on a dense LEO constellation at the lowest feasible orbital altitudes (currently primarily $\sim 550$ km and selectively $\sim 350$ km subject to coordination with NASA). In contrast, its main competitor, AST SpaceMobile relies on a less dense constellation at higher altitudes ($\sim 700$ km), supported by very large antenna arrays. Although both operators have so far relied primarily on IMT spectrum obtained through agreements with mobile network operators in different countries, they are now pursuing their own terrestrial and satellite spectrum. Other actors in the emerging DS2D ecosystem include Lynk Global, Skylo, and Viasat-Inmarsat. Key use cases comprise ubiquitous connectivity for IoT, supplementary coverage for terrestrial mobile networks, SOS alerts, emergency communications, and maritime and aeronautical connectivity.
@@ II-B Starlink’s Direct Satellite-to-Device (DS2D) Approach
In November 2024, the FCC granted partial approval for Starlink’s DS2D operations to initiate the SMS-only beta testing, but imposed several conditions to prevent harmful interference with terrestrial and satellite incumbents, including the requirement to stop transmission immediately upon any substantiated claim of interference, limitations on power flux density levels for aggregate out-of-band emissions (OOBE), and mandatory coordination procedures with incumbents. [^1] See FCC’s DA 24-1193
Subsequently, in March 2025, the FCC approved [^2] See FCC’s DA 25-197 a conditional waiver allowing Starlink to increase its OOBE by $9.4$ dB, despite formal objections from terrestrial operators, including AT&T and Verizon, and EchoStar. The waiver stipulated that Starlink must continue to adhere to strict OOBE limits and engage in active interference resolution with affected parties.
In the U.S., Starlink’s DS2D beta operations have relied exclusively on the PCS G Block, a $2\times 5$ MHz channel at 1910191019151915 MHz (uplink) and 1990199019951995 MHz (downlink), which is part of the IMT spectrum licensed to T-Mobile across the contiguous United States. [^3] See FCC’s IBFS File No.: SAT-LOA-20210511-00064 In September 2025, having concluded the beta phase, Starlink acquired Mobile Satellite Service (MSS) spectrum rights from Echostar in the AWS-4 band (2000200020202020 MHz uplink, 2180218022002200 MHz downlink, $2\times 20$ MHz) and IMT PCS H Block (1915191519201920 MHz uplink, 1995199520002000 MHz downlink, $2\times 5$ MHz). These additional allocations significantly expand the spectrum available for its Direct-to-Cell operations, particularly since the AWS-4 band overlaps with the 5G NR NTN band n252.
@@ II-C Related Work
Following initial deployments and the establishment of operational networks, a growing number of measurement studies have emerged, primarily focusing on Starlink’s fixed satellite broadband services  [1, 7, 8, 9]. These studies have primarily evaluated connectivity delivered through dedicated fixed user terminals (i.e., satellite “dishes”), with particular attention to metrics such as throughput and latency, a major historical limitation of traditional geostationary satellite systems.
@@ IV-C Analysis of DS2D Key Performance Indicators
In this Subsection, we present key metrics of Starlink’s DS2D RAN, namely RSRP, RSRQ and SINR. We examined the dataset for spatial patterns but did not identify any significant trends. In the temporal domain, we observed slight changes. Fig. 5 shows biweekly distributions (in the form of letter-value plots) of RSRP, RSRQ and SINR measurements from the first week of October 2024 (week 41) to the last week of July 2025 (week 32). This timeline captures key milestones of the service setup, including the authorization granted in March 2025 to increase OOBE by $9.4$ dB. The latter, due to transmitter non-linearities, would allow for an approximately $3.1$-dB increase in transmit power according to SpaceX declarations. [^5] See David Goldman’s (Space-X) letter in response to Merissa Velez (FCC) on September 13, 2024. Electronic Comment Filing System (ECFS) code 10914226871971/1 A detailed analysis shows a slight improvement in the distribution of the RAN metrics over the analyzed period, although the observed increment, in particular in RSRP, is smaller than the expected value.
@@ IV-D Performance assessment of future DS2D services
Building on the previous observations, we evaluate the potential performance of future DS2D data services. For this, we apply the modified Shannon capacity formulation (see Section III-C) using the available bandwidth and SINR measurements of the last three months, to ensure that the results reflect a more stable operational state of the network. Based on the SINR distributions, the mean and median values correspond to spectral efficiencies of $0.61$ and $0.52$ bps/Hz, respectively. Note that, given that high SINR values are rarely observed, this estimation is largely insensitive to the method used to compute $m$, and the influence of this parameter is negligible.
Since Starlink’s current DS2D operations are restricted to the PCS G Block ($2\times 5$ MHz), the derived spectral efficiency corresponds to an approximate downlink capacity of $3.1$ Mbps per beam. This value should be regarded as an upper-bound estimate of the throughput per connection, corresponding to the scenario of a single user occupying the full bandwidth of the beam. While such speeds may be sufficient to provide basic connectivity in remote or underserved regions, they remain well below the average downlink rates offered by terrestrial networks. This gap highlights the complementary nature of DS2D services in their current phase, consistent with their role as an SCS solution.
@@
First, as discussed in Subsection IV-C, we do not observe the expected $3.1$-dB increase in the RAN metrics that would be expected from applying the FCC’s March 2025 authorization; instead, only a slight improvement is reflected in the measurements, which, with the available evidence, it is not possible to determine whether this variation is attributable to the regulatory change. It therefore remains unclear whether transmissions are already occurring at the newly authorized power levels, although SpaceX’s filings suggest that this is the case (see footnote 4). Nevertheless, in the hypothetical assumption that the $9.4$-dB relaxation in the OOBE limit allows for a future increase in transmit power, and assuming that the increase in SINR would be similar (coverage-limited network), we estimate that the spectral efficiency would increase from the current baseline to $0.89$ bps/Hz (mean) and $0.83$ bps/Hz (median), raising the estimated average per-beam capacity from $3.1$ Mbps to $4.5$ Mbps.
Second, system performance may also be improved by lowering orbital altitudes. SpaceX has been authorized to selectively operate DS2D satellites in Very Low Earth Orbits (VLEO) at an altitude of 340340360360 km, subject to previous coordination with NASA. [^6] See FCC’s DA 24-1193 Assuming free-space propagation, relocating the entire constellation to lower altitudes would reduce path loss by $20\log_{10}(550/355)\approx 3.8$ dB. To comply with the OOBE constraints, this would require an estimated $1.3$-dB reduction in transmit power to maintain the power flux density levels for OOBE, resulting in a net gain of about $2.5$ dB in received signal power. Assuming again that the SINR values are increased in the same proportion, the spectral efficiency would raise from its baseline value to $0.83$ bps/Hz (mean) and $0.76$ bps/Hz (median). This corresponds to an approximate $1.4\times$ gain, yielding an estimated downlink capacity of $4.2$ Mbps per beam.
@@
In the longer term, satellite swarm antenna arrays show promising preliminary results in increasing the directivity of satellite beams [15], which allows for higher gains and improved spatial multiplexing. Achieving this would require the launch and operation of many more satellites, which might also contribute to increasing the overall throughput of DS2D networks and the effective capacity experienced by end users. Although the FCC denied Starlink’s request to authorize 30 00030\,000 DS2D-capable satellites, the company remains authorized to operate up to 7 5007\,500 such satellites, indicating substantial remaining potential for network densification. Furthermore, the aforementioned advancements in network capacity could be combined, potentially enabling greater per-beam throughput in the future.

### v8

@@ I Introduction
Following technological developments in fixed broadband connectivity, satellite-based cellular communications, known as Direct Satellite-to-Device (DS2D), have become a promising solution to provide seamless connectivity directly to standard unmodified smartphones. Starlink, the satellite communications division of SpaceX, has emerged as the global leader in DS2D communications by rapidly deploying a non-3GPP-NTN-compliant solution that leverages terrestrial International Mobile Telecommunications (IMT) spectrum. In partnership with T-Mobile and using its spectrum, Starlink has been the first company to launch large-scale beta testing of a Supplemental Coverage from Space (SCS) service. The system is initially limited to Short Message Service (SMS), with support for voice and data services planned by late 2025.
@@ II-A Technological Evolution and Emerging Ecosystem
Starlink has pioneered DS2D deployments with a strategy focused on a dense LEO constellation at the lowest feasible orbital altitudes (currently primarily $\sim 550$ km and selectively $\sim 350$ km subject to coordination with NASA). In contrast, its main competitor, AST SpaceMobile, relies on a less dense constellation at higher altitudes ($\sim 700$ km), supported by very large antenna arrays. Although both operators have so far relied primarily on IMT spectrum obtained through agreements with mobile network operators in different countries, they are now pursuing their own terrestrial and satellite spectrum. Other actors in the emerging DS2D ecosystem include Lynk Global, Skylo, and Viasat-Inmarsat. Key use cases comprise ubiquitous connectivity for IoT, supplementary coverage for terrestrial mobile networks, SOS alerts, emergency communications, and maritime and aeronautical connectivity.
@@ II-B Starlink’s Direct Satellite-to-Device (DS2D) Approach
In November 2024, the FCC granted partial approval for Starlink’s DS2D operations to initiate the SMS-only beta testing, but imposed several conditions to prevent harmful interference with terrestrial and satellite incumbents, including the requirement to stop transmission immediately upon any substantiated claim of interference, limitations on power flux density levels for aggregate out-of-band emissions (OOBE), and mandatory coordination procedures with incumbents. [^1] See FCC’s DA 24-1193.
Subsequently, in March 2025, the FCC approved [^2] See FCC’s DA 25-197. a conditional waiver allowing Starlink to increase its OOBE by $9.4$ dB, despite formal objections from terrestrial operators, including AT&T and Verizon, and EchoStar. The waiver stipulated that Starlink must continue to adhere to strict OOBE limits and engage in active interference resolution with affected parties.
In the U.S., Starlink’s DS2D beta operations have relied exclusively on the PCS G Block, a $2\times 5$ MHz channel at 1910191019151915 MHz (uplink) and 1990199019951995 MHz (downlink), which is part of the IMT spectrum licensed to T-Mobile across the contiguous United States. [^3] See FCC’s IBFS File No.: SAT-LOA-20210511-00064. In September 2025, having concluded the beta phase, Starlink acquired Mobile Satellite Service (MSS) spectrum rights from Echostar in the AWS-4 band (2000200020202020 MHz uplink, 2180218022002200 MHz downlink, $2\times 20$ MHz) and IMT PCS H Block (1915191519201920 MHz uplink, 1995199520002000 MHz downlink, $2\times 5$ MHz). These additional allocations significantly expand the spectrum available for its Direct-to-Cell operations, particularly since the AWS-4 band overlaps with the 5G NR NTN band n252.
@@ II-C Related Work
Following initial deployments and the establishment of operational networks, a growing number of measurement studies have emerged, focusing on Starlink’s fixed satellite broadband services  [1, 7, 8, 9]. These studies have primarily evaluated connectivity delivered through dedicated fixed user terminals (i.e., satellite “dishes”), with particular attention to metrics such as throughput and latency, a major historical limitation of traditional geostationary satellite systems.
@@ IV-C Analysis of DS2D Key Performance Indicators
In this Subsection, we present key metrics of Starlink’s DS2D RAN, namely RSRP, RSRQ and SINR. We examined the dataset for spatial patterns but did not identify any significant trends. In the temporal domain, we observed slight changes. Fig. 5 shows biweekly distributions (in the form of letter-value plots) of RSRP, RSRQ and SINR measurements from the first week of October 2024 (week 41) to the last week of July 2025 (week 32). This timeline captures key milestones of the service setup, including the authorization granted in March 2025 to increase OOBE by $9.4$ dB. The latter, due to transmitter non-linearities, would allow for an approximately $3.1$-dB increase in transmit power according to SpaceX declarations. [^5] See David Goldman’s (SpaceX) letter in response to Merissa Velez (FCC) on September 13, 2024. Electronic Comment Filing System (ECFS) code 10914226871971/1 A detailed analysis shows a slight improvement in the distribution of the RAN metrics over the measurement period, although the observed increment, in particular in RSRP, is smaller than the expected value.
@@ IV-D Performance assessment of future DS2D services
Building on the previous observations, we evaluate the potential performance of future DS2D data services. For this, we apply the modified Shannon capacity formulation (see Section III-C) using the available bandwidth and SINR measurements of the last three months, to ensure that the results reflect a more stable operational state of the network. Based on the SINR distributions, the mean and median spectral efficiency values are $0.61$ and $0.52$ bps/Hz, respectively. Note that, given that high SINR values are rarely observed, this estimation is largely insensitive to the method used to compute $m$, and the influence of this parameter is negligible.
Since Starlink’s current DS2D operations are restricted to the PCS G Block ($2\times 5$ MHz), the derived spectral efficiency corresponds to an approximate downlink capacity of $3.1$ Mbps per beam. This value should be regarded as an upper-bound estimate of the average throughput per connection, corresponding to the scenario of a single user occupying the full bandwidth of the beam. While such speeds may be sufficient to provide basic connectivity in remote or underserved regions, they remain well below the average downlink rates offered by terrestrial networks. This gap highlights the complementary nature of DS2D services in their current phase, consistent with their role as an SCS solution.
@@
First, as discussed in Subsection IV-C, we do not observe the $3.1$-dB increase in the RAN metrics that would be expected from applying the FCC’s March 2025 authorization; instead, only a slight improvement is reflected in the measurements. With the available evidence, it is not possible to determine whether this variation is attributable to the regulatory change. It therefore remains unclear whether transmissions are already occurring at the newly authorized power levels, although SpaceX’s filings suggest that this is the case (see footnote 4). Nevertheless, in the hypothetical assumption that the $9.4$-dB relaxation in the OOBE limit allows for a future increase in transmit power, and assuming that the increase in SINR would be similar (coverage-limited network), we estimate that the spectral efficiency would increase from the current baseline to $0.89$ bps/Hz (mean) and $0.83$ bps/Hz (median), raising the estimated average per-beam capacity from $3.1$ Mbps to $4.5$ Mbps.
Second, system performance may also be improved by lowering orbital altitudes. SpaceX has been authorized to selectively operate DS2D satellites in Very Low Earth Orbits (VLEO) at an altitude of 340340360360 km, subject to previous coordination with NASA. [^6] See FCC’s DA 24-1193 Assuming free-space propagation, relocating the entire constellation to lower altitudes would reduce path loss by $20\log_{10}(550/355)\approx 3.8$ dB. To comply with the OOBE constraints, this would require an estimated $1.3$-dB reduction in transmit power, resulting in a net gain of about $2.5$ dB in received signal power. Assuming again that the SINR values are increased in the same proportion, the spectral efficiency would raise from its baseline value to $0.83$ bps/Hz (mean) and $0.76$ bps/Hz (median). This corresponds to an approximate $1.4\times$ gain, yielding an estimated downlink capacity of $4.2$ Mbps per beam.
@@
In the longer term, satellite swarm antenna arrays show promising preliminary results in increasing the directivity of satellite beams [15], which allows for higher gains and improved spatial multiplexing. Achieving this would require the launch and operation of many more satellites, which might also contribute to increasing the overall throughput of DS2D networks and the effective capacity experienced by end users. Although the FCC denied Starlink’s request to authorize 30 00030\,000 DS2D-capable satellites, the company remains authorized to operate up to 7 5007\,500 such satellites, indicating substantial potential for network densification. Furthermore, the aforementioned advancements in network capacity could be combined, potentially enabling greater per-beam throughput in the future.